OpenStudy (anonymous):

Convert -450° from degrees to radians. Use the value of π found on a calculator and round answers to four decimal places, as needed.

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

how many degrees in a circle?

OpenStudy (anonymous):

you have to divide -450 by 180 correct?

OpenStudy (anonymous):

better to think of it as a conversion, then i think it's easier to keep straight... \(360^{\circ} = 2\pi \text{ radians}\) we use this to make a conversion factor. since you are given degrees and want to convert to radians, radians needs to be on top and degrees on the bottom:\[-450^{\circ}\cdot \frac{ 2 \pi }{ 360^{\circ} } = ? \text{ radians}\] just multiply and divide... the units cancel leaving radians

OpenStudy (anonymous):

or simplify then perform the operations... the result will be the same.